The consumption of electricity in a country from the past 9 years is studied by the power management department. The department has found that the increase in the consumption of electricity follows a linear model. The first year value of the consumption was 51753 (Gwh) and the consumption of the electricity in the 3rd year was 58128 (GWh). Keeping this information, Answer the following questions.

(a) Find the rate at which the electricity consumption changing with respect to time.

(b) Interpret the answer of part (a) in context of the question.

(c) Find an equation that connect the electricity consumption (y) and time (x).

(d) Find the electricity consumption in the 9th year by using the equation obtained in part (c).
1
Expert's answer
2020-06-12T12:01:53-0400

(a) The electricity consumption increases by "(58128 - 51753)\/(3 - 1) = 3187.5" per year.

(b) The answer of part (a) means that the average increase in electricity consumption is 3187.5 (Gwh) per year.

(c) The equation that connect the electricity consumption (y) and time (x) is:

y = 51753 + 3187.5(x - 1)

(d) The electricity consumption in the 9th year by using the equation obtained in part (c) is:

"y = 51753 + 3187.5\u00d7(9 - 1) = 77253."